OpenStudy (anonymous):

-(1+7)-6(-7-x)=36 how does one work this step by step

OpenStudy (anonymous):

What about -(1+7x)-6(-7-x)=36?

OpenStudy (anonymous):

x = 5

OpenStudy (anonymous):

Can you explain to me how to work it?

OpenStudy (anonymous):

-1 - 7x + 42 + 6x = 36 41 - x = 36 x = 5